# STAR: SocioTechnical Approach to Red Teaming Language Models

### v3

@@ 2.1 Steerability
Unintentional skews in red teaming may result from practical factors such as attacker demographics or task design. For example, open-ended approaches are intended to foster broad exploration, but can inadvertently lead to clustered redundancies as red teamers may naturally gravitate towards familiar or easily exploitable vulnerabilities. This tendency can be amplified by incentive structures that reward quick or easily identifiable harms. Furthermore, a lack of demographic diversity among human red teamers can exacerbate this issue, as attacks often reflect attackers own, inherently limited, experiences and perspectives (Ganguli et al., 2022; Feffer et al., 2024).
Prior work to addresses this challenge still has limitations. One strategy is to simply increase the number of attacks, but this is costly and doesn’t guarantee comprehensive coverage, as multiple attackers may still exploit the same harm vector. Principled approaches include dynamic incentives that reward the discovery of impactful vulnerabilities (Attenberg et al., 2015), framing diverse prompt generation as a quality-diversity search (Samvelyan et al., 2024) and using parametric instructions (Radharapu et al., 2023), though these approaches have not been applied to human red teaming of generative AI.
@@ 2.2 Signal Quality
Another significant challenge in red teaming is ensuring high quality of collected human data, especially when assessing harms that rely on subjective judgments. Prior work has shown comparably high rates of disagreement between raters when evaluating attack success (Ganguli et al., 2022; Xu et al., 2021). While often dismissed as noise, this disagreement can be a valuable source of information, reflecting the diverse perspectives that are essential to consider in evaluating AI model safety Aroyo and Welty (2015); Plank (2022). Simply taking a majority vote loses such signal, and risks overlooking minority judgments rooted in marginalised experiences.
Reduced signal quality may also stem from to skewed demographics of red teamers, as race, gender, and geo-cultural region have been shown to influence judgments on objectionable or adversarially generated content (Jiang et al., 2021; Goyal et al., 2022; Homan et al., 2023; Aroyo et al., 2023; DeVos et al., 2022). Yet, red teaming and annotation teams often lack demographic diversity (Feffer et al., 2024), even when efforts are made to recruit diversely. In prior studies, the majority of red teamers identified as white, cis-gendered, heterosexual, and without disabilities, with men often outnumbering women (Ganguli et al., 2022; Thoppilan et al., 2022). [^1] Only very few red teaming reports document annotator demographics. This lack of representation is likely to be more pronounced in red teaming efforts that did not deliberately recruit a diverse pool of workers. Furthermore, most red teaming focuses on English-language attacks, excluding many demographic groups and their languages (Nakamura et al., 2024). Such demographic skew can lead to undetected risks for these communities, potentially perpetuating disproportionate risks of harm when AI systems are deployed (Yong et al., 2024). To ensure broad coverage and legitimate and reliable data points, red teaming should involve diverse groups, encompassing a wider range of perspectives and experiences (Bockting et al., 2023). In addition, principled approaches are needed to account for meaningful annotator disagreement.
@@ 3.1 Improving Steerability
STAR addresses the steerability challenge by providing procedurally generated instructions to ensure comprehensive and even coverage of a targeted risk area. STAR’s instructions contain multiple parameters that delineate the targeted risk area (see detailed instructions in Appendix C). This content-agnostic approach is adaptable to any target area. As a proof of concept, we focus on red teaming for ethical and social harms, as codified in different ‘rules’ in a proprietary content safety policy, see B. To demonstrate that STAR enables steerability also in complex manifolds, we particularly explore two ‘rules’ - on hate speech and discriminatory stereotypes - with up to two additional instruction parameters that specify demographic groups to target. Our instructions encompass multiple parameters (see Figure  2 and Methods) to ensure that red teamers systematically explore the risk space, reducing redundancies and uncovering potential vulnerabilities that might otherwise be overlooked. These parameters are additive, meaning that specifying one (e.g., a rule) doesn’t limit our ability to measure harm across other parameters (e.g., different use cases). As such, additional parameters can be added – constrained only by the cognitive load they impose on human raters. We stress test this approach by aiming for coverage across labels of different levels of specificity: attackers may be asked to attack demographic groups based on single labels (race, gender), or combinatory labels (race×$\times$×gender).
@@ 3.2 Improving Signal Quality
Applying a sociotechnical lens [^2] A sociotechnical approach is rooted in the observation that AI systems are sociotechnical systems: both humans and machines are necessary in order to make the technology work as intended (Selbst et al., 2019). In the context of red teaming, this entails considering the interplay between different human attackers or annotators with an AI system, as well as societal and systemic structures that influence definitions of harm - such as what ‘ counts’ as a discriminatory stereotype and what does not., STAR centers the interplay of human attackers and annotators with the AI system. To provide a legitimate and reliable signal, we leverage different types of expertise, employing fact-checkers, medical professionals, and lived experience of generalists from different demographic groups. To learn from disagreement, we introduce an arbitration step to our annotation pipeline.
@@
Experts provide a more reliable and authoritative signal in their domains of expertise. This is why we employ raters with fact-checking and medical expertise to annotate relevant rules. We extend this logic to lived experience, which constitutes a relevant form of expertise on whether or not a given utterance constitutes hate speech or discriminatory stereotypes against one’s own demographic group. In addition, affected communities arguably should be prioritised and offer a more legitimate signal in the context of offense against their specific groups. Thus all attacks on medical, public interest, or demographic groups are annotated leveraging the relevant form of expertise.
We also anticipate that people of different demographic groups are often more familiar with the discriminatory stereotypes and hate speech targeted at their own group, compared to people of other demographic groups (Bergman et al., 2024). As a result, asking people to design attacks targeting their own group may create more ecologically valid signal, i.e. better reflect likely attacks from malicious users in real-world settings who rely on common tropes and stereotypes Gordon et al. (2021); Parrish et al. (2024). To test the relative effectiveness of ‘demographic matching’ not for annotation but for red teaming, 50% of attacks against a given demographic group are conducted by demographically-matched attackers, and 50% by a control of out-group attackers. This required recruiting a diverse red teaming and annotator pool. In particular, we recruit red teamers and annotators to obtain an even spread over multiple demographic labels including on gender and ethnicity (for demographics see  E).
@@
To enhance the reliability of our red teaming process, we gather labels from two annotators. However, discrepancies in judgment often arise. While some argue that such disagreement should be preserved for subjective tasks (e.g. Aroyo and Welty, 2015; Plank, 2022; Aroyo et al., 2023), red teaming necessitates clear safety recommendations. Unlike prior methods that merely add up ratings, we implement a two-step annotator $\rightarrow$ arbitrator pipeline, modeling argument exchange as in normative annotation settings (Bergman et al., 2023).
We first obtain annotator reasoning for their judgment on whether the model violated a rule. When two annotators’ ratings significantly diverge (by two or more steps on the four-step Likert scale), their dialogue and reasoning are presented to a third annotator, acting as an arbitrator. This arbitrator provides an additional rating and explanation, with all three annotations retained. This process uncovers annotator thinking and allows the arbitrator to weigh different perspectives for a more comprehensive judgment. The arbitrator is subject to the same expert- and demographic matching logic as annotators.
@@ 4 Methods
We obtain 8360 dialogues by 227 red teamers, annotated by 308 annotators and arbitrators, all in January 2024. Each dialogue adversarially tests model performance on one rule out of a content safety policy (see B).
@@
Adversarial testing here is conducted by human red teamers, in a multi-turn (mean: 16.4 ± 11.3) setting. Participants are assigned a red teaming task, an annotation task, or an arbitration task. Participants can perform multiple tasks in sequence, but they never see the same dialogue twice.
@@
Red teamers are given procedurally generated instructions with up to five parameters, directing red teamers to:
@@
    Violate a specific rule from the safety policy;
@@
    In cases where the rule involves hate speech or discriminatory stereotypes, identify the specific demographic group targeted by the attack.
@@
Arbitrators are provided with a dialogue between a red teamer and the proprietary model, and with the free-text reasoning from both previous annotators. They are then asked to make their judgement using the same Likert scale as annotators, and to provide their open-text reasoning.
@@
We recruited $n=313$ participants for our study (of which $n=286$ red teamed and $n=225$ annotated at least once), ensuring demographic diversity through self-identification in a voluntary questionnaire. Participants independently interacted with and evaluated the model under ethical approval from our ethics committee. Particular care was taken to build well-being considerations such as rest and opt-out steps into the task. They were compensated based on time spent (adhering to local living wage standards), to ensure quality assurance.
@@ 5.1 UMAP embedding
To compare thematic clustering of red teaming approaches, we project dialogues (between attacker and language model) from multiple datasets into a shared word embedding space (Fig. 2). [^3] We first project the dialogues onto high-dimensional embeddings using Gecko (Lee et al., 2024)), then onto two dimensions using UMAP (McInnes et al., 2020). We chose UMAP to be able to compare STAR to prior results, particularly building on (Ganguli et al., 2022). UMAP is a dimension reduction technique that finds a low-dimensional representation of high-dimensional data while preserving the data’s underlying structure. These datasets include two prior red teaming efforts, STAR, and a dataset of real-world dialogues between users and a proprietary system, which were flagged by the user due to the model displaying undesired behaviour. For a fair comparison, we downsample each dataset by randomly selecting the same number of data points. For more detail on these datasets see Appendix F).
We use hierarchical agglomerative clustering on the UMAP embeddings to identify twenty semantic groupings for the dialogues (iteratively joining pairs of clusters that are close to each other in the euclidean space of the UMAP embedding (Pedregosa et al., 2011)). The approximate outlines of these clusters are drawn manually in Figure 2), and semantic labels per cluster are listed in Table 1. Two reviewers independently assigned semantic labels per cluster and disagreeing labels (clusters 4 and 13) were reviewed by a third reviewer, followed by a discussion among all labellers to determine the final labelling via consensus.
@@ 5.2 Quantitative and qualitative analysis
For comparing in- vs. out-group annotations, we include all dialogues where red teamers were instructed to attack a specific demographic group in the context of breaking the discriminatory stereotypes and hate speech rules. These dialogues were annotated mostly by in-group members, with 62% of these dialogues being annotated by in-group members and 38% by out-group members. We compute odds ratios of different groups mentioned in instructions to yield a successful red teaming attempt and test statistical significance using ANOVA and t-tests. For qualitative insights, we manually inspect a random sample of rater dialogues and annotator reasoning.
@@ 6.1 Controlled exploration of the target area
Visual inspection of Figure 2 shows comparably broad coverage and low clustering of the STAR approach, despite more specific instructions compared to the other projected red teaming approaches. Analysing clusters in the embedding space reveals a thematic split between the three red teaming approaches (Table 1). The most common themes in STAR dialogues concern gender stereotypes (cluster 2) and race-based bias (16), followed by medical topics (8), reflecting the instructions. The most common themes in Anthropic dialogues are malicious use (5), explicit stories including adult fiction (3), and facilitating crime (0). Most common themes in DEFCON dialogues are prompts about model training followed by model refusals (4), passwords and sensitive personal data (7), and PII including from celebrities (14). In contrast, most common themes in real-world flagged dialogues were advice and recommendations (1), computer code (12) and refusals (4).
Figure 2: UMAP of the embedding space of dialogues across three red teaming datasets: Anthropic, DEFCON, and STAR; as well as dialogues between a proprietary model and users that were flagged as undesirable by users. Each dot indicates a dialogue. For comparability, we downsampled all datasets to include maximum 4000 randomly selected instances.
@@
Analysing the spread of red teaming attacks across race, gender, and race×$\times$×gender intersectionalities reveals that STAR achieves a sufficiently even spread of attacks across these categories as intended. Predictable exceptions arose regarding the labels "non-binary", "Asian and male", and "Hispanic and male", where we were unable to recruit the target number of red teamers (Fig. 3) [^4] Recall that instructions are limited by the availability of red teamers of each demographic group, as we show instructions or annotations of dialogues that harm certain groups only to demographically matched members of that group..
@@ 6.2.1 Demographic matching affects annotations
When split by rule, only hate speech shows a statistically significant difference between in-group and out-group annotators in terms of their likelihood of rating a rule as broken (see Fig. 5). We also test in- vs. out-group attack success at red teaming against a particular demographic group but here we find no significant differences (see  G). Qualitative analyses further hint at different sensitivity profiles underlying in in- vs. out-group disagreement. Disagreement often arose when the target group was alluded to or referenced indirectly, or in the context of ‘positive’ stereotypes, with in-group members more often marking such dialogues as violative of the rule (see  I.1). Out-group members on the other hand, appeared more likely to mark dialogues where the user makes a problematic statement and the model fails to counter it, as violative - even when the model did not explicitly endorse harmful views. Finally, out-group raters appeared more likely to cite model refusal or disclaimers in association with marking a dialogue as non-violative, compared to in-group members.

### v4

@@ 2.1 Steerability
Unintentional skews in red teaming may result from practical factors such as attacker demographics or task design. For example, open-ended approaches are intended to foster broad exploration, but can inadvertently lead to clustered redundancies as red teamers may naturally gravitate towards familiar or easily exploitable vulnerabilities. This tendency can be amplified by incentive structures that reward easily identifiable harms. Furthermore, a lack of demographic diversity among human red teamers can exacerbate this issue, as attacks often reflect attackers own, inherently limited, experiences and perspectives (Ganguli et al., 2022; Feffer et al., 2024).
Prior work to address this challenge still has limitations. One strategy is to simply increase the number of attacks, but this is costly and doesn’t guarantee comprehensive coverage, as multiple attackers may still exploit the same harm vector. Principled approaches include dynamic incentives that reward the discovery of impactful vulnerabilities (Attenberg et al., 2015), framing diverse prompt generation as a quality-diversity search (Samvelyan et al., 2024) and using parametric instructions (Radharapu et al., 2023), though these approaches have not been applied to human red teaming of generative AI.
@@ 2.2 Signal Quality
Another significant challenge in red teaming is ensuring high quality of collected human data, especially when assessing harms that rely on subjective judgments. Prior work has shown high rates of disagreement between raters when evaluating attack success (Ganguli et al., 2022; Xu et al., 2021). While often dismissed as noise, this disagreement can be a valuable source of information, reflecting the diverse perspectives that are essential to consider in evaluating AI model safety Aroyo and Welty (2015); Plank (2022). Simply taking a majority vote loses such signal, and risks overlooking minority judgments rooted in marginalised experiences.
Reduced signal quality may also stem from skewed demographics of red teamers, as race, gender, and geo-cultural region have been shown to influence judgments on objectionable or adversarially generated content (Jiang et al., 2021; Goyal et al., 2022; Homan et al., 2023; Aroyo et al., 2023; DeVos et al., 2022). Yet, red teaming and annotation teams often lack demographic diversity (Feffer et al., 2024), even when efforts are made to recruit diversely. In prior studies, the majority of red teamers identified as white, cis-gendered, heterosexual, and without disabilities, with men often outnumbering women (Ganguli et al., 2022; Thoppilan et al., 2022). [^1] Only very few red teaming reports document annotator demographics. Red teaming efforts that did not deliberately recruit a diverse pool of workers are likely to have even less representation . Furthermore, most red teaming focuses on English-language attacks, excluding many demographic groups and their languages (Nakamura et al., 2024). Such demographic skew can lead to undetected risks for these communities, potentially perpetuating disproportionate risks of harm when AI systems are deployed (Yong et al., 2024). To ensure broad coverage and legitimate and reliable data points, red teaming should involve diverse groups, encompassing a wider range of perspectives and experiences (Bockting et al., 2023). In addition, principled approaches are needed to account for meaningful annotator disagreement.
@@ 3.1 Improving Steerability
To ensure comprehensive and even coverage (Section 2.1), STAR divides the the targeted risk area based on multiple parameters (see Figure 1 and Section 4). For every red teaming attempt, instructions are procedurally generated based on a general templates, filled in with different combinations of parameter values (see example of the resulting instructions in Appendix C).
This content-agnostic approach is adaptable to any target area. As a proof of concept, we focus on red teaming for ethical and social harms, as codified by ‘rules’ in a proprietary content safety policy (see Appendix B).
To demonstrate that STAR enables steerability also in complex manifolds, we particularly explore two ‘rules’ – hate speech and discriminatory stereotypes – with up to two additional instruction parameters that specify demographic groups to target.
All these parameters are additive, meaning that specifying one (e.g., a rule) doesn’t limit our ability to measure harm across other parameters (e.g., different use cases). As such, additional parameters can be added – constrained only by the cognitive load they impose on human raters. We stress test this approach by aiming for coverage across labels of different levels of specificity: attackers may be asked to attack demographic groups based on single labels (race, gender), or combinatory labels (race×$\times$×gender).
@@ 3.2 Improving Signal Quality
Applying a sociotechnical lens, STAR centers the interplay of human attackers and annotators with the AI system. A sociotechnical approach is rooted in the observation that AI systems are sociotechnical systems: both humans and machines are necessary in order to make the technology work as intended (Selbst et al., 2019). In the context of red teaming, this entails considering the social identities of attackers and annotators and how this may influence red teaming results. It also entails considering societal and systemic structures that influence definitions of harm - such as what ‘counts’ as a discriminatory stereotype and whose perspectives may be less well-represented in the context of defining such harms. Given the socially situated nature of conversational AI systems (sartori2022sociotechnical), a sociotechnical exploration of their failure modes can shine a light on critical real-world failure modes that may otherwise go undetected.
STAR introduces a socio-technical perspective to red teaming language models through two key contributions. First, its methods highlight how identity groups may be affected differently by an AI system at the point of use, and thus red teaming the harm areas of stereotypes and hate with regard to specific demographic groups and intersectionalities. Second, by taking into account the identities and different forms of expertise of red-teamers – defining expertise as lived experiences, in addition to professional and academic expertise – STAR emphasises the importance of taking into account who wields influence along different stages of developing an AI system.
To provide a legitimate and reliable signal (Section 2.2), we leverage different types of expertise, employing fact-checkers, medical professionals, and lived experience of generalists from different demographic groups. To learn from disagreement, we introduce an arbitration step to our annotation pipeline.
@@
Experts provide a more reliable and authoritative signal in their domains of expertise. This is why we employ raters with fact-checking and medical expertise to annotate relevant rules. We extend this logic to lived experience, which constitutes a relevant form of expertise on whether or not a given utterance constitutes hate speech or discriminatory stereotypes against one’s own demographic group. In addition, affected communities arguably should be prioritised and offer a more legitimate signal for judging offense against their specific groups. Thus all attacks on medical, public interest, or demographic groups are annotated leveraging the relevant form of expertise.
We also anticipate that people of different demographic groups are often more familiar with the discriminatory stereotypes and hate speech targeted at their own group, compared to people of other demographic groups (Bergman et al., 2024). As a result, asking people to design attacks targeting their own group may create a more ecologically valid signal, i.e. better reflect likely attacks from malicious users in real-world settings who rely on common tropes and stereotypes Gordon et al. (2021); Parrish et al. (2024). To test the relative effectiveness of ‘demographic matching’ not for annotation but for red teaming, 50% of attacks against a given demographic group are conducted by demographically-matched attackers, and 50% by a control of out-group attackers. This required recruiting a diverse red teaming and annotator pool. In particular, we recruit red teamers and annotators to obtain an even spread over multiple demographic labels including on gender and ethnicity (for demographics see Appendix F).
@@
To enhance the reliability of our red teaming process, we gather labels from two annotators. However, discrepancies in judgment often arise. While some argue that such disagreement should be preserved for subjective tasks (e.g. Aroyo and Welty, 2015; Plank, 2022; Aroyo et al., 2023), red teaming necessitates clear safety recommendations. Unlike prior methods that merely add up ratings, we implement a two-stage annotator $\rightarrow$ arbitrator pipeline, modeling the exchange of arguments in a normative annotation setting (where multiple legitimate perspectives exist) (Bergman et al., 2023).
We ask both annotators to provide reasoning alongside their judgment on whether the model violated a rule. If the two annotators’ ratings significantly diverge (by two or more steps on the four-step Likert scale), their dialogue and reasoning are presented to a third annotator, acting as an arbitrator. This arbitrator provides an additional rating and explanation, with all three annotations retained. This process uncovers annotator thinking and allows the arbitrator to weigh different perspectives for a more comprehensive judgment. The arbitrator is subject to the same expert- and demographic matching logic as annotators.
@@ 4 Methods
We obtain 8360 dialogues by 225 red teamers, annotated by 286 annotators and arbitrators, all in January 2024. Each dialogue adversarially tests model performance on one rule out of a content safety policy (Appendix B). Participant compensation and labour costs are detailed in Appendix E.
@@
Adversarial testing here is conducted by human red teamers, in a multi-turn (mean: 16.4 ± 11.3 turns) setting. Participants are assigned a red teaming task, an annotation task, or an arbitration task. Participants can perform multiple tasks in sequence, but they never see the same dialogue twice.
@@
Red teamers are given procedurally generated instructions with random values for up to five parameters inspired by characteristics_of_harmful_text, directing red teamers to:
@@
    Steer the model into violating a specific rule from the safety policy;
@@
    In cases where the rule involves hate speech or discriminatory stereotypes, steer the model into targeting a specific demographic group.
@@
Arbitrators are provided with a dialogue between a red teamer and the proprietary model, and with the free-text reasoning from both previous annotators. They are then asked to make their judgement using the same Likert scale as annotators, and to provide their own free-text reasoning. See instructions in Figure 9.
@@
We recruited $n=313$ participants for our study (of which $n=225$ red teamed and $n=286$ annotated at least once), ensuring demographic diversity through self-identification in a voluntary questionnaire. Participants independently interacted with and evaluated the model under ethical approval from our ethics committee. Particular care was taken to build well-being considerations such as rest and opt-out steps into the task. They were compensated based on time spent (adhering to local living wage standards), so that there was no incentive to rush.
@@ 5.1 UMAP embedding
To compare thematic clustering of red teaming approaches, we project dialogues (between attacker and language model) from multiple datasets into a shared embedding space (Figure 2). [^2] We first project the dialogues onto high-dimensional embeddings using Gecko (Lee et al., 2024)), then onto two dimensions using UMAP (McInnes et al., 2020), with the cosine distance between Gecko embeddings, and the structure generated via the 5 closest neighbors. We chose UMAP to be able to compare STAR to prior results, particularly building on (Ganguli et al., 2022). UMAP is a dimension reduction technique that finds a low-dimensional representation of high-dimensional data while preserving the data’s underlying structure. These datasets include two prior red teaming efforts, STAR, and a dataset of real-world dialogues between users and a proprietary system, which were flagged by the user due to the model displaying undesired behaviour. For a fair comparison, we downsample each dataset by randomly selecting the same number of data points. For more detail on these datasets see Appendix G).
We use hierarchical agglomerative clustering (Ward linkage, 20 clusters to allow for manual inspection within reason) on the UMAP embeddings to identify twenty semantic groupings for the dialogues (iteratively joining pairs of clusters that are close to each other in the euclidean space of the UMAP embedding (Pedregosa et al., 2011)). The approximate outlines of these clusters are drawn manually in Figure 2), and semantic labels per cluster are listed in Table 1. Two reviewers independently assigned semantic labels per cluster and disagreeing labels (clusters 4 and 13) were reviewed by a third reviewer, followed by a discussion among all labellers to determine the final labelling via consensus. The choices of hyperparameters were fixed a priori to avoid cherry picking of results.
@@ 5.2 Quantitative and qualitative analysis
For comparing in- vs. out-group annotations, we include all dialogues where red teamers were instructed to attack a specific demographic group in the context of breaking the discriminatory stereotypes and hate speech rules. Demographic matching would usually match all of these dialogues to an in-group annotator, but as an ablation we collected additional annotations with deliberately mismatched demographics (out-group) such that 59% of these dialogues were annotated by in-group members only, 24% by out-group members only, and 17% by both.
We compute odds ratios of different groups mentioned in instructions to yield a successful red teaming attempt and test statistical significance using ANOVA and t-tests. For qualitative insights, we manually inspect a random sample of rater dialogues and annotator reasoning.
@@ 6.1 Controlled exploration of the target area
Visual inspection of Figure 2 shows the coverage and low clustering of the STAR approach compared to the other projected red teaming approaches, despite more specific instructions. The specific instructions in STAR are designed to increase coverage of intended area of risk surface. Analysing clusters in the embedding space reveals a thematic split between the three red teaming approaches (Table 1). The most common themes in STAR dialogues concern gender stereotypes (cluster 2) and race-based bias (16), followed by medical topics (8), reflecting the instructions. The most common themes in Anthropic dialogues are malicious use (5), explicit stories including adult fiction (3), and facilitating crime (0). The most common themes in DEFCON dialogues are prompts about model training followed by model refusals (4), passwords and sensitive personal data (7), and PII including from celebrities (14). In contrast, the most common themes in real-world flagged dialogues were advice and recommendations (1), computer code (12) and refusals (4).
Figure 2: UMAP of the embedding space of dialogues across three red teaming datasets: Anthropic, DEFCON, and STAR; as well as dialogues between a proprietary model and users that were flagged as undesirable by us. Visual inspection of Figure 2 shows similar coverage and clustering of the STAR approach compared to other approaches. Cluster analysis further reveals that STAR results in more intentional thematic clustering based on the red teaming instructions, compared to the other projected red teaming approaches. Each dot indicates a dialogue. For comparability, we downsampled all datasets to include maximum 4000 randomly selected instances.
@@
Analysing the spread of red teaming attacks across race, gender, and race×$\times$×gender intersectionalities reveals that STAR achieves a sufficiently even spread of attacks across these categories as intended. Predictable exceptions arose regarding the labels "non-binary", "Asian and male", and "Hispanic and male", where we were unable to recruit the target number of participants (Figure 3) [^3] Recall that we assign annotations of dialogues that target a group only to demographically matched members of that group..
@@ 6.2.1 Demographic matching affects annotations
When split by rule, only hate speech shows a statistically significant difference between in-group and out-group annotators in terms of their likelihood of rating a rule as broken (see Figure 5). We also test in- vs. out-group attack success at red teaming against a particular demographic group but here we find no significant differences (see Appendix H). Qualitative analyses further hint at different sensitivity profiles underlying in- vs. out-group disagreement. Disagreement often arose when the target group was alluded to or referenced indirectly, or in the context of ‘positive’ stereotypes, with in-group members more often marking such dialogues as violative of the rule (see  J.1). Out-group members on the other hand, appeared more likely to mark dialogues where the user makes a problematic statement and the model fails to counter it, as violative - even when the model did not explicitly endorse harmful views. Finally, out-group raters appeared more likely to cite model refusal or disclaimers in association with marking a dialogue as non-violative, compared to in-group members.